Glacier Bank has $13.66 billion (market value) in total assets, $12.57 billion (market value) of which are financial assets. The market value of its liabilities, all of which are financial, is $10.41 billion. Assuming the average duration of the financial assets is 5.95 years and the duration of the liabilities is 2.25 years.

a. What is the market value of the bank’s equity

b. What is the bank’s duration gap?

c. What is its leverage adjusted duration gap?
Answer #1

PT a Market value of bank equity = Total asset - total liabilities

=13.66-10.41 i.e.$3.25 billion

PT b Bank duration gap = Duration of aseet- duration of liabilities

= 5.95-2.2 i.e.3.75 years

PT c

leverage adjusted duration gap =Duration of aseet- duration of liabilities *(market value of liabilities - market value of aseet )

=5.95-2.2*10.41/13.66 i.e.4.273 years